the radius of yhe planet saturn is 5.87×10^7m and it's mass is 5.86×10^26kg. find the density of its material by g/m find it's surface area by m2 (4πr)
1
Expert's answer
2019-08-26T10:30:49-0400

The density of its material:


"\rho=\frac{m}{V}=\frac{3m}{4\pi r^3}"

"\rho=\frac{3(5.86\cdot 10^{26})}{4\pi (5.87\cdot 10^7)^3}"

"\rho=692\frac{kg}{m^3}"

It's surface area:


"A=4\pi r^2=4\pi (5.87\cdot 10^7)^2=4.33\cdot 10^{16}\ m^2"